6 unit multipliers to convert 500 cubic feet to cubic cm?
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

500 ft³ = 14158423.356 cm³
http://www.metric-conversions.org/volume/cubic-feet-to-cubic-centimeters.htm

Answer by stanbon(75887)

6 unit multipliers to convert 500 cubic feet to cubic cm?
(500 ft^3)(12 in/1 ft)^3*(1 cm/0.39 in)^3 = (500*12^3*1^3)/(1^3*0.39)^3 (cm)^3
------
= (864000/(0.0593190) (cm)^3
---
= 14565316.34 cm^3
===========================
Cheers,
Stan H.